bob has 60 coins consisting of quarters and dimes. The coins combined value is $9.45. Find out how many are quarters and how many are dimes
Answer by mananth(16949)

quarters ------------ x
dimes ------------ y
1 x + 1 y = 60 .............1
25 x + 10 y = 945 .............2

multiply (1)by -10
Multiply (2) by 1
-10 x + -10 y = -600
25 x + 10 y = 945

Add the two equations

15 x=345

/15

x=23 quarters

plug value of x in (1)

x+y=60
23+y=60
y = 60 + -23
y = 37 dimes